OpenStudy (anonymous):

I also need help on how to find the prime factorization of 540. How do you write it in exponential form

OpenStudy (kainui):

Well if a number ends in an even number like 0,2,4,6, or 8 then you know it's divisible by 2. If a number ends in 0 or 5 then it must also be divisible by 5. So you can start by dividing out one number at a time. For instance if I give you: 20 then divide out 2 or 5 first, it doesn't matter, let's try 2. 20/2=10 So we can say 20=2*10 Since 10 isn't prime let's keep going, let's try 5 now 10/5=2 So now we have 20=2*2*5 which are all prime numbers. But notice that we have 2 twice! That's the same as 2 squared. So we write it like this: \[20=2^2*5\] I hope that example helps, try it out and if you get stuck I'll help you more.

OpenStudy (anonymous):

Thank you kainui but Im still stuck.

OpenStudy (kainui):

Where? I can help you further if you give me more information. I don't know what else to say unless you tell me what's not working out.